OpenStudy (anonymous):

Precalculus trigonometry identities : Cot^3t/csct=cos t (csc^2 t-1)

OpenStudy (anonymous):

\[ \cot t = \frac{\cos t}{\sin t} \]

OpenStudy (anonymous):

\[ \csc t = \frac 1{\sin t} \]

OpenStudy (anonymous):

What do you do after ?